In Lopez Company, total material costs are $36,000, and total conversion costs are $54,000. Equivalent units of production are materials 10,000 and conversion costs 12,000.

Compute the unit costs for materials, conversion costs, and total manufacturing costs. (Round answers to 2 decimal places, e.g. 2.25.)

Materials cost per unit  

 

Conversion cost per unit  

 

Total manufacturing cost per unit  
Answer #1
Concepts and reason

Cost of Material: The cost incurred in the purchase of the materials required to manufacture a product is considered as the cost of material. This cost is influenced by the volume of production.

Fundamentals

Manufacturing Cost: The costs incurred in processing the raw materials in order to convert them into finished products is called manufacturing costs. They basically consist of the cost of supplies, direct expenses and like.

Conversion Cost: The cost incurred in the conversion of the raw materials into finished goods with the help of direct labor and other manufacturing overheads. This cost includes the direct labor cost and other overheads incurred in the manufacturing of the product.

Direct cost: It is the cost directly related to the manufacture of the product. It includes the direct material, direct labor and direct expenses incurred during production. The direct cost is influenced by the volume of production.


‎Overheads Cost: Also known as operating costs, the expenses that are incurred on all the operations or activities of a company other than those costs that are directly expensed on the production processes are called overheads cost.

Equivalent Units of Production: The units which are almost complete which reflect the amount of work done with the available resources. The completed units and the almost completed units are expressed in terms of fully completed units.

Compute the unit cost of material.

Unitcostofmaterial=TotalCostofMaterialEquivalentunitsofproduction=$36,00010,000units=$3.60perunit.\begin{array}{c}\\{\rm{Unit cost of material}} = \frac{{{\rm{Total Cost of Material}}}}{{{\rm{Equivalent units of production}}}}\\\\ = \frac{{\$ 36,000}}{{10,000\,{\rm{units}}}}\\\\ = \$ 3.60\,{\rm{per unit}}{\rm{.}}\\\end{array}

Compute the per unit conversion cost.

Unitcostofconversioncost=TotalConversionCostEquivalentunitsofproduction=$54,00012,000units=$4.50perunit\begin{array}{c}\\{\rm{Unit cost of conversion cost}} = \frac{{{\rm{Total Conversion Cost}}}}{{{\rm{Equivalent units of production}}}}\\\\ = \frac{{\$ 54,000}}{{12,000\,{\rm{units}}}}\\\\ = \$ 4.50\,{\rm{per}}\,{\rm{unit}}\\\end{array}

Compute the total manufacturing cost per unit.

(TotalManufacturingCostPerUnit)=Unitmaterialcost+UnitConversionCost=$3.60+$4.50=$8.10\begin{array}{c}\\\left( \begin{array}{l}\\{\rm{Total Manufacturing }}\\\\{\rm{Cost Per Unit}}\\\end{array} \right) = {\rm{Unit material cost}} + {\rm{Unit Conversion Cost}}\\\\ = \$ 3.60 + \$ 4.50\\\\ = \$ 8.10\\\end{array}

Ans:

The total cost of material 36,000 is divided by the equivalent units of production as material as the basis 10,000. The resultant unit cost of the material is $3.60.

The total cost of conversion cost 54,000 is divided by the equivalent units of production as conversion cost as the basis 12,000. The resultant conversion cost per unit is $4.50.

The total manufacturing costs is computed by adding the material cost per unit $3.60 along with the conversion cost per unit $4.50. The resultant total manufacturing cost per unit is $8.10.